Consider a system consisting of two concentric sol
Answer #1

n= 105/cm outer solenoid
n= 125/cm for inner solenoid.

A) B= uo*n* i -------> 4pi *10^-7* 1050*1.25/ 10^-4= 16.5 mT
B) B= B inner- B outer ----> 34.1-16.5= 17.6
B inner = 4pi *10^-7*1250*2.17/10^-4=34.08 left